In a certain bimetallic strip, the brass strip is 0.100% longer than the steel strip at a temperature of 287°C. At what temperature do the two strips have the same length? Coefficients of linear expansion for steel α = 12.0 × 10−6 K−1 and for brass α = 19.0 × 10−6 K−1

1
Expert's answer
2022-05-16T16:06:51-0400

"\frac{1+\alpha_1∆t}{1+\alpha_2∆t}=1.001,"

"t_2=t_1+\frac{0.001}{\alpha_2-1.001\alpha_1}=144°C."